# S2C2A: A Flexible Task Space Planning and Control Strategy for Modular Soft Robot Arms

### v1

@@ III-C Configuration Estimation
To collect accurate MSRA states and configurations, most methods use an external sensing systems like the optical tracking system shown in Figure 2-(A) or electromagnetic tracking systems in [60] for accurate feedback control. However, such systems have strict requirements on optical or electromagnetic environments and are unsuitable for real applications like agriculture and domestic environments. In this work, we aim to perform MSRA state control only by employing internal sensing feedback. Encoders in the motors are employed in this work, and some other internal sensors such as IMU[61] and flex sensor[62] may also be utilized for replacing the encoder and providing internal sensing feedback in future work.
@@
The encoder in the motor can measure cable displacement and serves as an internal module configuration sensor. Following [63], we estimate module orientation based on the cable displacement. Three cables $c_{1},c_{2},c_{3}$ are placed in each module, and $O$ represents the plane center as shown in Figure 2-(C). In the initial state, the length of each cable and the module is $l_{0}$. During the motion, we define the cable lengths as $l_{1},l_{2},l_{3}$ and the cable displacement $a_{i}=l_{i}-l_{0}$, where $i=1,2,3$. Considering the low axial direction compressibility of each module, we assume that the module length does not change, and the displacements follow:
@@
where $r_{p}$ represents the distance between the plane center and the cable position, as shown in Figure 2-(C). Then, the configuration $(o_{x},o_{y},o_{z})$ can be denoted as
@@
In this case, we can estimate the module configuration $(o_{x},o_{y},o_{z})$ based on encoder feedback $a_{1},a_{2},a_{3}$.
III-D Data Collection
Using the experiment devices mentioned in Sec. III-A, we collected 9000 data points, as depicted in Figure 3-(A) and (B), and the working space length of the MSRA end on the x, y, and z axis are about 0.99, 0.99, and 0.49 m. Instead of a purely random motor babbling strategy, we actuate the MSRA in different patterns, as illustrated in Figure 3-(C). All the modules are actuated with the same random action sequence $a_{I}$ in the first one-third period. Then, in the second one-third period, the first and second modules are actuated with the same random action sequence $a_{I}$ while the end module is actuated by a different action sequence $a_{III}$. In the final one-third period, all these modules are actuated by different random action sequences $a_{I},a_{II},a_{III}$. This data collection approach specifically for MSRAs is proposed in [16] and has demonstrated that it can explore a larger working space in the 3D space and varied module configuration combinations.
@@ III-D Data Collection
Figure 3: (A) Some collected samples. Samples with $x>0$ and $y<0$ are hidden for visualization. The positions of the base module end, middle module end, and end module end are depicted as blue, green, and red dots. (B) Data collection strategy specifically for MSRAs.
Figure 4: The middle module bending angles estimated by the optical tracking system (ground truth, red) and encoders in motors (red).
@@
We compare the real module configurations collected by the optical tracking system (ground truth) and the configurations estimated by encoder feedback following Equation 3. The estimation errors of the base, middle, and end modules are 7.02, 8.07, and 12.77, respectively. An example of a comparison of the ground truth and estimated module bending angles during dynamic motion is shown in Figure 4. The increase in error along the module sequence could be attributed to gravity, hysteresis, and friction between the cables and modules. In this work, instead of using the optical tracking system as the configuration sensor, we aim to achieve state control even under the inaccurate configuration estimation. Such an implementation endows the MSRA with a tracking ability that relies only on internal sensing feedback instead of costly external sensing systems like optical and electromagnetic tracking systems. Internal sensing correction may be included in future work.
IV Planning and Control Methods
[…]
IV-A Configuration Planning - S2C
@@
Our planning strategy is inspired by model predictive control (MPC), which determines action based on a forward model and an optimization problem. By adjusting the cost function in the optimization problem, diverse tasks can be performed. Considering the nonlinearity, hysteresis, and modularity of MSRAs, it is challenging to build an accurate physical model. Hence we leverage a biLSTM network $NN_{C2S}$ as the forward model estimating state $(\hat{p^{\prime}},\hat{o^{\prime}})$ based on module configuration $C$, as shown in Figure5-(A). All the data, including robot states, configurations, and actions, are rescaled to $[-1,1]$ for NN training. In this paper, we apply $I,II,III$ to represent the module number and $1\sim 5$ to represent the time step.
@@
In order to fulfill various tasks, a series of losses is integrated into the cost function. To achieve position-tracking and orientation-tracking tasks, we introduce the losses
$\begin{split}L_{p}&={\|p_{d}^{\prime}-\hat{p^{\prime}}\|}_{2},\\ L_{o}&={\|o_{d}^{\prime}-\hat{o^{\prime}}\|}_{2},\end{split}$ (4)
where $p_{d}^{\prime},o_{d}^{\prime}$ are the desired state position and orientation from the target states.
@@
where $d={\|p_{ob}^{\prime}-\hat{p^{\prime}}\|}_{2}$ refers to the distance between the estimated position $\hat{p^{\prime}}$ and the obstacle center $p_{ob}^{\prime}$. $r$ defines the obstacle avoidance threshold according to the obstacle risk level.
@@
$\begin{split}{\min_{C}}\ \mu_{p}L_{p}+&\mu_{o}L_{o}+\mu_{ob}L_{ob}+\mu_{d}L_{d},\\ {s.t.}\ (\hat{p^{\prime}},\hat{o^{\prime}})&=NN_{C2S}(C)\\ \end{split}$ (7)
where $\mu_{p},\mu_{o},\mu_{ob},\mu_{d}$ represent the weights of position control, orientation control, obstacle avoidance, and configuration change. $NN_{C2S}$ serves as a forward model estimating the robot state $(\hat{p^{\prime}},\hat{o^{\prime}})$ from the module configurations $C$, and the module configurations $C$ are optimized to minimize the cost function.
Compared to other NN implementations in soft robots which only target a single task, such a planning strategy is adaptive to various tasks. Moreover, this planning strategy can be implemented both offline and online, depending on the task requirements.
IV-B Configuration Control - C2A
Following [16], we utilize a biLSTM $NN_{C2A^{\prime}}$ and a ’sum and range’ layer to compose $NN_{C2A}$ as the configuration controller for MSRAs, as shown in Figure 5-(B). The controller can be represented as
@@ IV-B Configuration Control - C2A
The actions of each module $\hat{A_{m,t}}\in R^{3}$ should follow the constraint in Equation 1 and in the range $[-1,1]$. Hence, one layer follows the biLSTM
$\begin{split}\hat{A_{m,t}}=\frac{3tanh(\hat{A_{m,t}^{\prime}})-\sum\tanh(\hat{A_{m,t}^{\prime}})}{4},\end{split}$ (10)
@@
Considering the uniform deformation of our cable-driven robots, a physical controller based on the CC model can also be applied for configuration control. To decide the action $\hat{a_{1}},\hat{a_{2}},\hat{a_{3}}$ based on the desired configuration $(o_{dx},o_{dy},o_{dz})$, we first calculate the desired bending angle $\phi_{d}$ and neutral surface direction $\theta_{d}$ by
$\begin{split}\theta_{d}&=atan2(o_{dy},o_{dx})+\frac{\pi}{2},\\ \phi_{d}&=atan2(\frac{o_{dy}}{sin(\theta_{d}-\frac{\pi}{2})},o_{dz}).\\ \end{split}$ (11)
Then, we decide the actions by
$\begin{split}\hat{a_{1}}&=-\phi_{d}r_{p}sin(\theta_{d}),\\ \hat{a_{2}}&=\phi_{d}r_{p}sin(\frac{2\pi}{3}-\theta_{d}),\\ \hat{a_{3}}&=\phi_{d}r_{p}sin(\theta_{d}-\frac{\pi}{3}).\\ \end{split}$ (12)
Compared to the physical controller, the data-driven controller $NN_{C2A}$ is a more general controller available for various soft robots, like the pneumatic MSRA in [16] and the cable-driven MSRA in this work. However, this physical controller is only available to the soft robots following the CC assumption. Different modules require specific physical models and the targeted controller, but physical approaches have higher explainability than the black-box NN controller. Overall, the S2C2A strategy can implement different configuration controllers based on the MSRA properties and task requirements.
Utilizing the dataset collected in Section III-D, we train $NN_{C2S}$ and $NN_{C2A}$ for configuration planning and control, and the NN hyperparameters are shown in Table I. Of note, the configuration applied in NN training is only from internal sensors, not the optical system. $NN_{C2A}$ contains fewer layers and a smaller hidden state size to decrease the online computation burden.
TABLE I: NN hyperparameters
@@
optimizer Adma Adam
average estimation error 1.36$\pm$1.211.07$\pm$1.89%
To minimize the cost functions in Equation 7 and generate the target configuration online or offline, we employ Pytorch[64], and the optimizer is Adam. The learning rate is 0.02, and the optimization iteration is 10.
V Experimental Results
V-A Position Control
We propose a position trajectory in the task space, as shown in the red lines in Figure 6, and employ our approach to achieve position control. Following the target trajectory, the MSRA will first bend along the x-axis and rotate to move along a circle. We include the losses $L_{p}$ and $L_{d}$ for position tracking and compare the performance of $NN_{C2A}$ and the physical CC controller. In this task, $u_{p}=1$ and $u_{d}=0.5$. The errors of trajectories using NN controller and CC controller are shown in Table II. The MSRA trajectories and robot motions are shown in Figure 6 and 7. All the experiment videos can be found in the Supplementary Video.
@@ V-A Position Control
TABLE II: Errors and standard derivations in position control
C2A position error
$NN$ 2.8$\pm$1.0 cm
$CC$ 3.7$\pm$1.4 cm

### v2

@@ II-C Configuration Estimation
To collect accurate MSRA states and configurations, most works use external sensing systems like the optical tracking system shown in Fig. 2(A) or electromagnetic tracking systems in [22] for accurate feedback control. However, such systems have strict requirements on optical or electromagnetic environments and are unsuitable for real applications like agriculture and domestic environments. In this work, we aim to perform MSRA state control by only employing rough internal sensing feedback. Encoders in the motors are employed in this work, and some other internal sensors such as IMU[12] and flex sensor[23] may also be utilized for correcting the encoder and providing internal sensing feedback in future work.
@@
The encoder in the motor can measure cable displacement and serve as an internal sensor. We reconstruct the configuration based on the PCC model because it is one of the simplest soft robot models and produces a clear mapping from the cable displacement to the module configuration. Following [24], we estimate module orientation based on the cable displacement. Three cables $c_{1},c_{2},c_{3}$ are placed in each module, and $O$ represents the plane center as shown in Figure 2(C). In the initial state, the length of each cable and the module is $l_{0}$. During the motion, we define the cable lengths as $l_{1},l_{2},l_{3}$ and the cable displacement $a_{i}=l_{i}-l_{0}$, where $i=1,2,3$. Considering the low axial direction compressibility of each module, we assume that the module length does not change, and the displacements follow:
@@
where $r_{p}$ represents the distance between the plane center and the cable position, as shown in Fig. 2(C). Then, the configuration $[o_{x},o_{y},o_{z}]$ can be denoted as
@@
In this case, we can estimate the module configuration $[o_{x},o_{y},o_{z}]$ based on encoder feedback $a_{1},a_{2},a_{3}$.
Of note, the encoder cannot provide accurate configuration estimations due to cable slack, friction, and a possible mismatch between the real robots and the PCC model. In this work, we do not aim to address this error. Instead, we bridge the mapping between this internal configuration space and the task and actuation space, as shown in Fig. 1(C). Using this kind of rough and inaccurate configuration, our strategy can achieve various MSRA state control tasks thanks to our NN approaches. In future work, we may only utilize raw feedback instead of estimation based on PCC.
II-D Data Collection
@@
Figure 3: (A) Some collected samples. Samples with $x>0$ and $y<0$ are hidden for visualization. The positions of the base module end, middle module end, and end module end are depicted as blue, green, and red dots. The yellow dot represents the MSRA base. (B) Collected samples exhibited on the x-z plane. (C) Data collection strategy specifically for MSRAs.
When collecting data from MSRA, the purely motor babbling strategy applied in most neural network works [10] may result in data samples confined to a limited space. To explore a larger area and reach the boundaries of the MSRA working space, all modules must bend in the same direction. However, this scenario is rare when module actuations are independent and random by the purely motor babbling strategy, especially with the increase of the module number. In this case, we utilize a data collection strategy specifically for MSRA following [7], as shown in Fig. 3(B). All the modules are actuated with the same random action sequence $a_{I}$ in the first one-third period. Then, in the second one-third period, the base and middle modules are actuated with the same random action sequence $a_{I}$ while the end module is actuated by a different action sequence $a_{III}$. In the final one-third period, all these modules are actuated by different random action sequences $a_{I},a_{II},a_{III}$. Leveraging our data collection strategy, the MSRA can reach the working space edge in the beginning while exploring different configuration combinations after that.
Using the experiment devices mentioned in Sec. II-A, we collected 9000 data points, as depicted in Fig. 3(A), and the working space length of the MSRA end on the x, y, and z axis are about 0.92 m, 0.92 m, and 0.39 m. Meanwhile, the samples collected by the pure motor babbling strategy are constrained in a $0.87m\times 0.87m\times 0.34m$ space, which is smaller than ours. This data collection approach specifically for MSRAs has demonstrated that it can explore a larger working space and variegated module configuration combinations both on our robot and in [7].
@@
Based on the collected data, we compare the real module configurations collected by the optical tracking system (ground truth) and the configurations estimated by encoder feedback following Equation 3. The average estimation errors of the base, middle, and end modules are 2.6, 5.1, 8.9, and the maximal errors are 8.8, 19.6, 36.3, respectively. By taking the middle module as an example, the comparison of the ground truth and estimated module bending angles during dynamic motion is shown in Fig. 4. The increase in error along the module sequence could be attributed to gravity, cable slack, and friction between the cables and modules.
Figure 4: The middle module bending angles estimated by the optical tracking system (ground truth, red) and encoders in motors (green).
In this work, instead of using the optical tracking system as the configuration sensor or adjusting the estimation error, we focus on achieving state control even under the inaccurate configuration estimation, as shown in Fig. 1(C). Such an implementation endows the MSRA with a tracking ability that relies only on internal sensing feedback instead of costly external sensing systems like optical and electromagnetic tracking systems. Internal configuration correction may be included in future work.
Figure 5: (A) Space sequence biLSTM network $NN_{C2S}$. Each biLSTM unit takes configurations $C_{*}$ of one module as input to estimate the robot states $\hat{S_{*}}$. (B) Configuration Controller $NN_{C2A}$. A biLSTM network $NN_{C2A^{\prime}}$ decides the action $\hat{A_{*,t}}$ based on the previous actions $A_{*,t-1\sim t-5}$, target and previous states $C_{d*,t+1},C_{*,t\sim t-3}$, and module labels $n_{*}$. The ’sum and range’ layer is applied to guarantee the actuation sum and range.
III Planning and Control Methods
III-A Configuration Planning - S2C
@@
Our planning strategy is inspired by model predictive control (MPC), which determines action based on a forward model and an optimization problem. By adjusting the cost function in the optimization problem, diverse tasks can be performed. Considering the nonlinearity, hysteresis, and modularity of MSRAs, it is challenging to build an accurate physical model. Hence we leverage a biLSTM network $NN_{C2S}$ as the forward model estimating state $[\hat{p^{\prime}}\ \hat{o^{\prime}}]^{T}$ based on module configuration $C$, as shown in Fig. 5(A). All the data, including robot states, configurations, and actions, are rescaled to $[-1,1]$ for NN training.
@@
In order to fulfill various tasks, a series of losses is integrated into the cost function. To achieve position tracking and orientation tracking tasks, we introduce the losses
$\begin{split}L_{p}&={\|p^{\prime}_{d}-\hat{p^{\prime}}\|}_{2},\\ L_{o}&={\|o^{\prime}_{d}-\hat{o^{\prime}}\|}_{2},\end{split}$ (4)
where $p^{\prime}_{d},o^{\prime}_{d}$ are the desired position and orientation.
@@
where $d={\|p^{\prime}_{ob}-\hat{p^{\prime}}\|}_{2}$ refers to the distance between the estimated position $\hat{p^{\prime}}$ and the obstacle center $p^{\prime}_{ob}$. $r$ defines the obstacle avoidance threshold according to the obstacle risk level and obstacle size. In future work, we may update the obstacle center and avoidance threshold online for adjusting to the changing and moving obstacles.
@@
$\begin{split}{\min_{C}}\ \mu_{p}L_{p}+&\mu_{o}L_{o}+\mu_{ob}L_{ob}+\mu_{d}L_{d},\\ {s.t.}\ [\hat{p^{\prime}}\ \hat{o^{\prime}}]^{T}&=NN_{C2S}(C)\\ \end{split}$ (7)
where $\mu_{p},\mu_{o},\mu_{ob},\mu_{d}$ represent the weights of position control, orientation control, obstacle avoidance, and configuration change. $NN_{C2S}$ serves as a forward model estimating the robot state $[\hat{p^{\prime}}\ \hat{o^{\prime}}]^{T}$ from the module configurations $C$, and the module configurations $C$ are optimized to minimize the cost function. The loss weights can be adjusted according to task requirements, like the task propriety order; those applied in the following experiments have been decided after a trial and error process.
Compared to other optimization works using physical models mentioned in Sec. I, our data-driven strategy can reflect the real motion features of the MSRA. Compared to other NN implementations in soft robots, which only target a single task, such a planning strategy is adaptive to various tasks. Moreover, this planning strategy can be implemented both offline and online, depending on the task requirements.
Before the planning, $NN_{C2S}$ is trained using the dataset in Sec. II-D. To minimize the cost functions in Equation 7 and generate the target configuration online or offline, the $NN_{C2S}$ serves as a forward model, and its parameters are frozen during optimization. We employ Pytorch[25] for planning, and the optimizer is Adam. The learning rate is 0.02, and the optimization iteration is 10.
III-B Configuration Control - C2A
Following [7], we utilize a biLSTM $NN_{C2A^{\prime}}$ and a ’sum and range’ layer to compose $NN_{C2A}$ as the configuration controller for MSRAs, as shown in Fig. 5(B). The controller can be represented as
@@ III-B Configuration Control - C2A
Considering the low axial compressibility of our module, the actions of each module $\hat{A_{m,t}}\in R^{3}$ should follow the constraint $\sum_{i=1}^{3}\hat{A_{i,t}}=0$ to maintain the central axis length invariant following Equation 1. Also, they should be in the range $[-1,1]$ for NN training. Hence, one layer follows the biLSTM
$\begin{split}\hat{A_{m,t}}=\frac{3\tanh(\hat{A_{m,t}^{\prime}})-\sum_{i=1}^{3}\tanh(\hat{A_{i,t}^{\prime}})}{4},\end{split}$ (10)
@@
Overall, as shown in Fig. 5(B), the target configuration $C_{dm,t+1}$, previous configurations $C_{m,t}\sim C_{m,t-3}$, previous actions $A_{m,t-1}\sim A_{m,t-5}$, and module label $n_{m}$ are formatted into a 1-D vector and fed into the biLSTM unit $m$ as input, where $m=1,2,3$ represent the module number. The output of the unit $m$ is the estimated actuation $\hat{A_{m,t}}$ to reach the target configuration $C_{dm,t+1}$ considering the previous module configurations, actuations, and hidden states from the other units. BiLSTM is chosen as the configuration controller because it outperforms some other neural network controllers. The details of the biLSTM controller and the comparison with the other neural network controllers can be found in our previous work [7].
Utilizing the dataset collected in Section II-D, we train $NN_{C2S}$ and $NN_{C2A}$ for configuration planning and control, and the NN hyperparameters are shown in Table II. Of note, the configuration applied in NN training is only from internal sensors, not the optical system. The early stopping strategy is applied to avoid overfitting. These hyperparameters are chosen after trial and error, and $NN_{C2A}$ contains fewer layers and a smaller hidden state size to decrease the online computation burden.
TABLE II: NN hyperparameters
@@
optimizer Adam Adam
average estimation error 1.93$\pm$1.870.91$\pm$1.90%
IV Experimental Results
In this section, we demonstrate the versatility of our planning and control strategy in various tasks. First, we validate some of our strategies with primary experiments in simulation, including the data collection strategy and the robustness to sensing and actuation coupling in Sec. IV-A. Then, we compare our approach with our previous planning and controller strategy applied to our robot in [2]. The comparison results in position and orientation control are shown in Sec. IV-B and IV-C. Then, we validate our approach in some high-level tasks, like position constraint in Sec. IV-D, obstacle avoidance in Sec. IV-E, and online interaction in Sec. IV-F.
IV-A Simulation Validation
We apply a soft robot simulator based on the pseudo-rigid model as reported in [26], as shown in Fig. 6(B). A three-module MSRA is built in this simulator, and the length of each module is 0.2m. To validate the generalization of our strategies, four cables, rather than three cables in the real robots, are simulated in the robot, and the actuation value is force, not cable displacement.
Figure 6: (A) The dataset collected by our strategy in the simulation. The green, orange, and blue dots represent the end positions of the base, middle, and end modules. (B) The target MSRA end trajectory (blue) and the MSRA motion. Each module end is represented by one black dot, and its orientation is represented by a black vector. (C) The independent (left) and coupled (right) actuation diagram.
IV-A1 Data Collection Strategy Comparison
In Sec. II-D and our previous work [7], we have proven that our data collection strategy produces a larger working space than the common motor babbling strategy in MSRA. In this subsection, we also test the influence on control accuracy. First, we employ our strategy, including data collection strategy in Fig. 1(C), S2C planning, and C2A control, in simulation. 9000 samples are collected, as shown in Fig. 6(A). The error of this work is shown in Table III. The MSRA motion is shown in Fig. 6(B) and Fig. 7. The low error demonstrates that our strategy can achieve position control in MSRA.
Then, we apply the dataset collected by the common motor babbling strategy for neural network training, including $NN_{C2S}$ and $NN_{C2A}$. These two networks are applied for position control and named ’Motor babbling’ in Table III. The lower error of our strategy demonstrates that the dataset collected by our strategy produces a larger working space and contributes to accurate control. A dataset covering a larger working space provides richer configuration combinations, such as those reaching the edge, and enhances the neural network’s generalization ability to represent MSRA motion.
@@ IV-A1 Data Collection Strategy Comparison
TABLE III: Average errors and standard deviations comparison in position control in simulation
Approach Errors (cm)
This work $\bf{1.5\pm 0.3}$
Motor babbling 1.9$\pm$0.7
LSTM 3.3$\pm$1.7
Sensing coupling 1.7$\pm$0.3
Actuation coupling 1.6$\pm$0.5
Figure 7: The target (red) and measured motion of MSRA ends on the x, y, and z axes applying our control strategy (blue), motor babbling strategy (green), LSTM (light blue), under sensing coupling (purple) and actuation coupling (yellow).
IV-A2 Data-driven Approach Comparison
We compare our strategy with another data-driven approach widely applied in soft robot control, which is LSTM [10, 27]. In the existing works, the previous robot states and control target (end position in this task) are the LSTM input, while the LSTM decides the next actuation. The error of LSTM is higher than that of this work, as shown in Table III.
The NN applications in soft robotics are motivated by the specific challenges of the hardware. Multilayer perceptron is introduced in soft robotics to address the nonlinearity. Then, RNN targeted at sequence problems is widely applied due to the delayed motion. Following the same principle, when meeting challenges about modularity in MSRA, we utilize a novel NN, biLSTM, in planning and control. Unlike existing data-driven controllers, which apply target position as input to decide actuation, we utilize biLSTM as the forward model in optimization planning to cope with the complex state-to-configuration mapping caused by modularity.
IV-A3 Sensing Coupling Robustness
Module coupling is one of the most significant challenges in MSRA control. Some sensors in MSRA, like the encoders, pass through the whole MSRA and are affected by the module coupling. In this case, we built a sensing coupling model, denoted as
$\begin{split}\begin{bmatrix}\hat{C_{I}}\\ \hat{C_{II}}\\ \hat{C_{III}}\\ \end{bmatrix}=\begin{bmatrix}0.85&0.10&0.05\\ 0.20&0.70&0.10\\ 0.15&0.30&0.55\\ \end{bmatrix}\begin{bmatrix}{C_{I}}\\ {C_{II}}\\ {C_{III}}\\ \end{bmatrix},\end{split}$ (11)
where $\hat{C_{*}}$ represent the inaccurate configuration suffered from sensing coupling and ${C_{*}}$ represent the accurate configuration. We train our strategy with the inaccurate configuration and achieve position control, resulting in a similar error to ’Ours’ as shown in Table III. Similar errors demonstrate the robustness of our strategy against sensing coupling.
IV-A4 Actuation Coupling Robustness
In addition to sensing coupling, some actuation units, like cables, also pass through the whole MSRA, like the old version of our real MSRA mentioned in [2]. To test the robustness against actuation coupling, we couple the cables in simulation, as shown in Fig. 6(C). Then, we collect a dataset with this kind of MSRA and train our strategy with the new dataset. The error is shown in Table III, demonstrating that our strategy shows robustness against actuation coupling.
IV-B Position Control
We propose a position trajectory in the task space, as shown in the red lines in Fig. 8 and the white dotted trajectory in Fig. 9, and employ our biLSTM approach and PCC approach in [2] to achieve position control. Following the target trajectory, the MSRA will first bend along the x-axis and rotate to move along a circle. We include the losses $L_{p}$ and $L_{d}$ for position tracking and compare the performance of the PCC strategy [2] and our biLSTM strategy. In this task, $u_{p}=1$ and $u_{d}=0.5$. The errors of trajectories using the PCC and our strategies are shown in Table IV. The MSRA trajectories and robot motions are shown in Fig. 8 and 9. All the experiment videos can be found in the Supplementary Video.
TABLE IV: Average errors and standard deviations comparison in position and orientation control
task position error orientation error (z) orientation error (x)
p-PCC[2] 2.5$\pm$0.8 cm / /
p-ours $\bm{2.2\pm 1.1}$cm / /
40-PCC[2] 3.3$\pm$1.0 cm 2.1$\pm$1.4 /
40-ours $\bm{2.9\pm 0.8}$cm $\bm{2.1\pm 1.3}^{\circ}$ /
50-PCC[2] 3.9$\pm$1.4 cm 4.5$\pm$2.2 /
50-ours $\bm{2.3\pm 0.7}$cm $\bm{2.8\pm 1.4}^{\circ}$ /
60-PCC[2] 4.5$\pm$1.0 cm 7.5$\pm$3.7 /
60-ours $\bm{2.2\pm 1.1}$cm $\bm{3.5\pm 2.9}^{\circ}$ /
0-PCC[2] 2.9$\pm$0.8 cm 4.7$\pm$2.2 /
0-ours $\bm{2.6\pm 1.0}$cm $\bm{3.3\pm 1.0}^{\circ}$ /
50a20-PCC[2] 4.8$\pm$1.4 cm 3.9$\pm$2.3 6.8$\pm$3.4
50a20-ours $\bm{3.1\pm 1.6}$cm $\bm{2.4\pm 1.5}^{\circ}$ $\bm{3.6\pm 2.7}^{\circ}$
50c20-PCC[2] 3.8$\pm$1.1 cm 5.9$\pm$2.0 4.4$\pm$2.2
50c20-ours $\bm{3.1\pm 1.2}$cm $\bm{3.0\pm 2.3}^{\circ}$ $\bm{3.5\pm 3.1}^{\circ}$
Average-PCC[2] 3.7$\pm$1.1 cm 4.1$\pm$2.3 5.6$\pm$2.8
Average-ours $\bm{2.6\pm 1.1}$cm $\bm{2.9\pm 1.7}^{\circ}$ $\bm{3.6\pm 2.9}^{\circ}$